### v1

# Seeded Poisson Factorization: Leveraging domain knowledge to fit topic models

Abstract
Topic models are widely used for discovering latent thematic structures in large text corpora, yet traditional unsupervised methods often struggle to align with predefined conceptual domains. This paper introduces Seeded Poisson Factorization (SPF), a novel approach that extends the Poisson Factorization framework by incorporating domain knowledge through seed words. SPF enables a more interpretable and structured topic discovery by modifying the prior distribution of topic-specific term intensities, assigning higher initial rates to predefined seed words. The model is estimated using variational inference with stochastic gradient optimization, ensuring scalability to large datasets.
We apply SPF to an Amazon customer feedback dataset, leveraging predefined product categories as guiding structures. Our evaluation demonstrates that SPF achieves superior classification performance compared to alternative guided topic models, particularly in terms of computational efficiency and predictive performance. Furthermore, robustness checks highlight SPF’s ability to adaptively balance domain knowledge and data-driven topic discovery, even in cases of imperfect seed word selection. These results establish SPF as a powerful and scalable alternative for integrating expert knowledge into topic modeling, enhancing both interpretability and efficiency in real-world applications.
Keywords: Poisson factorization, topic model, variational inference, customer feedback
1 Introduction
Inferring latent structures in text data is a fundamental challenge in natural language processing and its application in a wide range of fields of research such as political science, social science and economics. Due to the unstructured nature of text data, text analysis poses distinct challenges compared to the analysis of other types of data that are commonly used in empirical research (see, e.g., Kelly et al.,, 2021). Topic modeling provides a widely used framework for discovering hidden thematic structures within text corpora, offering insights into the distribution of topics across documents and the association between words and topics. Among the available topic modeling approaches, in particular Latent Dirichlet Allocation (LDA; Blei et al.,, 2003) and its extensions (see, e.g., Lafferty and Blei,, 2005; Roberts et al.,, 2014; Eshima et al.,, 2024) have been widely studied and applied across various domains (see, e.g., Bagozzi and Berliner,, 2018; Barbera et al.,, 2019; Thorsrud,, 2020; Davis and Tabrizi,, 2021; Munro and Ng,, 2022; Zimmermann et al.,, 2024). However, alternative topic modeling frameworks, such as Poisson Factorization (PF), provide distinct advantages by leveraging a Poisson likelihood rather than a multinomial distribution.
PF has been shown to provide improved scalability and computational efficiency (see, e.g., Canny,, 2004; Gopalan et al.,, 2014, 2015). Unlike LDA, which relies on Dirichlet priors for topic distributions, PF leverages the Poisson distribution. PF factorizes the document-term matrices into non-negative latent components that naturally accommodate sparsity, making it well suited for large-scale text corpora (Gopalan et al.,, 2015). In the case of PF topic models, these latent components represent topic intensities over words $\bm{\beta}$ (referred to as topical content or topic-term intensities), and document intensities over topics $\bm{\theta}$ (referred to as topical prevalence or document-topic intensities). The topical content refers to what is being discussed, while the topical prevalence indicates how much it is being discussed. PF naturally promotes sparsity in the latent factors and can handle large datasets efficiently due to its inherent properties and the use of variational inference techniques (see, e.g., Vafa et al.,, 2020; Vávra et al., 2024a, ; Vávra et al., 2024b, ).
Despite the success of topic models in uncovering latent themes in textual data, traditional methods are often limited by their purely unsupervised nature. In many applications, researchers and practitioners require models that align with predefined conceptual domains or that allow for targeted analysis based on domain knowledge (see, e.g., Eshima et al.,, 2024, for an application in political science). Conventional topic models typically require extensive post-processing, such as the manual labeling of topics and the merging of semantically similar ones, to extract meaningful insights (see, e.g., Morstatter and Liu,, 2016; Eshima et al.,, 2024). To address this limitation, guided and semi-supervised topic models have been proposed, which incorporate human input a-priori to fit a topic model to learn topics of specific interest to a user (see, e.g., Jagarlamudi et al.,, 2012; Watanabe and Zhou,, 2022; Eshima et al.,, 2024). These approaches typically influence the topic-term distributions, allowing for more interpretable and controlled topic discovery. Recent advances in guided topic modeling have primarily focused on extending LDA-based models (Watanabe and Zhou,, 2022; Eshima et al.,, 2024; Watanabe and Baturo,, 2024), yet a corresponding extension within the PF framework has remained unexplored.
Figure 1: Architecture of the Seeded Poisson Factorization (SPF) topic model.
The paper at hand contributes to the literature as follows. Firstly, we introduce Seeded Poisson Factorization (SPF), a novel topic model that integrates domain knowledge into the PF framework through the inclusion of seed words. As shown in Figure 1, SPF extends standard PF by decomposing topic-term intensities into a neutral component and a seeded component informed by predefined seed words. This structured decomposition enables the model to incorporate prior knowledge while preserving the flexibility of PF to learn latent topic distributions from data. In this way, SPF learns specific topics of interest and adaptively adjusts for potential seed word mis-specifications by controlling the contribution of seeded components, ensuring robustness in applications where domain knowledge may be incomplete or imperfect. Secondly, in contrast to prior guided topic models, which predominantly rely on Markov Chain Monte Carlo (MCMC) inference (see, e.g., Eshima et al.,, 2024; Watanabe and Baturo,, 2024), SPF employs variational inference (VI) for scalable parameter estimation. VI formulates posterior inference as an optimization problem, significantly reducing computational costs compared to traditional sampling-based methods (Ranganath et al.,, 2014; Blei et al.,, 2017). We empirically demonstrate that SPF not only achieves competitive predictive performance but also exhibits substantial computational efficiency, making it particularly suitable for large-scale text corpora.
Topic models are used for automatic text analysis in a wide range of fields of research, including for example the analysis of speech data in political science (Vávra et al., 2024b, ) or open-ended survey responses in the social sciences (Roberts et al.,, 2014). In the following we focus on yet another area of application: automatic analysis of consumer feedback. Consumer feedback provides valuable insights into customer preferences, sentiment, and emerging trends (see, e.g., Khan and Jeong,, 2016; Filieri et al.,, 2018; Davis and Tabrizi,, 2021; Aghakhani et al.,, 2021; Biswas et al.,, 2022; Zhou et al.,, 2024). Given the growing volume of online reviews and their impact on decision-making, accurately categorizing and summarizing such feedback remains a central challenge in computational social science and business analytics. We thus make use of a publicly available Amazon customer review dataset to illustrate and evaluate the performance of SPF. In particular, we assess SPF’s performance to extract meaningful topics when seed words are supplied to characterize the underlying product categories of products discussed in the customer reviews. In addition we investigate how well SPF infers the product category of the product discussed in a consumer review by employing a Naive Bayes classifier on the inferred topic intensities. We also compare the predictive performance as well as the computational efficiency of SPF to competing recently proposed guided topic models with readily available software implementations, i.e., KeyATM (Eshima et al.,, 2024) and SeededLDA (Watanabe and Baturo,, 2024). Finally, we conduct a series of robustness checks to examine the sensitivity of SPF to variations in seed word quality, model specification and corpus characteristics. The results confirm that SPF provides excellent performance across diverse experimental conditions.
By introducing SPF, we contribute to the methodological literature on topic modeling by extending Poisson Factorization with domain-informed priors, providing an alternative to existing LDA-based guided topic models. Our results demonstrate that SPF enhances both interpretability and computational efficiency of topic models, offering a scalable solution for researchers and practitioners seeking structured topic discovery in large text corpora.
The rest of the paper is structured as follows. In Section 2 we describe our generative model. Section 3 outlines model inference. Section 4 provides empirical results by applying SPF to customer feedback data from Amazon. Section 5 concludes.
2 The Seeded Poisson Factorization model
Based on the bag-of-words assumption (see, e.g., Eshima et al.,, 2024) the data are summarized in a Document-Term-Matrix (DTM), $\mathbb{Y}$. This matrix has the dimension number of documents $D$ times number of unique terms (words) in the data $V$, where each row corresponds to a single document $d=1,\ldots,D$, and each column represents a specific term $v=1,\ldots,V$ from vocabulary $\mathscr{V}$. A single entry $y_{dv}$ contains the frequency count of term $v$ in document $d$, such that $y_{dv}\geq 0$. Poisson factorization topic models assume that the observed word frequencies are generated independently from a Poisson distribution. In our Seeded Poisson Factorization (SPF) model, we decompose the Poisson rates into a linear combination of document-topic intensities $\bm{\theta}$ and topic-term intensities $\bm{\beta}$ over latent topic dimension $K$ for every frequency count $y_{dv}$:
$y_{dv}\sim\mathsf{Pois}\left(\sum_{k=1}^{K}\theta_{dk}\beta_{kv}\right).$ (1)
Document-topic intensities form a tall matrix $\bm{\theta}=(\bm{\theta}_{d})_{d=1}^{D}=(\theta_{dk})_{d,k=1}^{D,K}$, while topic-term intensities form a wide matrix $\bm{\beta}=(\bm{\beta}_{k})_{k=1}^{K}=(\beta_{kv})_{k,v=1}^{K,V}$. The number of topics $K$ needs to be a-priori specified. Both intensity matrices consist of positive elements.
In our SPF model we include the prior knowledge available to researchers which suggests specific terms to be relevant for certain topics in the following way. We seed the topics by inflating the prior mean of topic-term intensities for seeded words. Let $\mathscr{V}_{k}\subset\mathscr{V}$ be the set of seed words for topic $k=1,\ldots,K$ of size $V_{k}=|\mathscr{V}_{k}|$. In practice, we expect only few seed words per topic, $V_{k}\ll V$ and denote by $\mathscr{S}=\bigcup_{k=1}^{K}\mathscr{S}_{k}$, $\mathscr{S}_{k}=\{(k,v),v\in\mathscr{V}_{k}\}$, the set of all seed words. We allow for $\mathscr{V}_{k}=\emptyset$, $V_{k}=0$, in which case the topic is not a-priori seeded. Note that in case $V_{k}=0$ for all $k=1,\ldots,K$, the SPF model reduces to a standard PF model.
For the topic-term intensities we assume that they can be decomposed into a component present for all terms and a component specific to seed words, i.e., $\beta_{kv}=\beta_{kv}^{\star}+\widetilde{\beta}_{kv}$ where $\widetilde{\beta}_{kv}>0$ for seeds and $\widetilde{\beta}_{kv}=0$ otherwise. Both components are given a gamma prior:
$\beta_{kv}^{\star}\sim\Gamma(a,b)\quad\text{and}\quad\widetilde{\beta}_{kv}\left\{\begin{aligned} &\sim\Gamma(c,d)&&\;\text{for}\;(k,v)\in\mathscr{S},\\ &=0&&\;\text{otherwise},\end{aligned}\right.$ (2)
with $a,b,c,d>0$ and $c\gg a$. In case $b=d$, this implies that $\beta_{kv}$ also has a gamma prior $\Gamma(a+c,b)$. All document-topic intensities $\theta_{dk}$ are given a gamma prior
$\theta_{dk}\sim\Gamma(e,f).$ (3)
To obtain the empirical results in Section 4, we set $a=b=e=f=0.3$, which according to Gopalan et al., (2015) results in sparse representations of the document-topic and topic-term intensities. To emphasize the relevance of the pre-defined seed words we set $c=1.0$ and $d=0.3$. Finally, the generative process is captured in plate notation in Figure 2.
Figure 2: Directed graphical representation of the SPF model. Shaded nodes are observed, transparent nodes are latent variables, double circles indicate deterministic transformations of parent nodes and points are fixed parameters.
3 Inference
3.1 Variational inference
Given the DTM $\mathbb{Y}$, we infer the document-topic intensities and the topic-term intensities based on approximating the posterior distribution over the model’s latent variables $p(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}}\,|\,\mathbb{Y})$. We use Variational Inference (VI) methods to fit an approximate posterior distribution (see, e.g., Blei et al.,, 2017). VI frames the inference as an optimization problem. The key steps of VI consist of selecting a parametric family of variational distributions $\mathscr{Q}=\{q_{\bm{\phi}},\bm{\phi}\in\bm{\Phi}\}$ and determining the parameter $\bm{\phi}^{*}\in\bm{\Phi}$ minimizing the Kullback-Leibler divergence ($\mathsf{KL}$) of the variational distribution from the true posterior
$q_{\bm{\phi}^{*}}(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})=\underset{q_{\bm{\phi}}\in\mathscr{Q}}{\operatorname{\mathrm{arg}\,\mathrm{min}\;}}\mathsf{KL}\left(q_{\bm{\phi}}(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})\,\left\|\,p(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}}\,|\,\mathbb{Y})\right.\right).$
This $\mathsf{KL}$ optimization problem is equivalent to maximizing the evidence lower bound ($\mathsf{ELBO}$):
$\mathsf{ELBO}(\bm{\phi})=\mathbb{E}_{{q}_{\bm{\phi}}}\left[\log p(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})+\log p(\mathbb{Y}\,|\,\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})-\log q_{\bm{\phi}}(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})\right]$ (4)
or minimizing the negative $\mathsf{ELBO}$ (Jordan et al.,, 1998; Bishop,, 2006). Equation 4 sums the expectation of the log-likelihood and the log-prior and the entropy of the variational family.
In the mean-field approach, the variational family $q_{\bm{\phi}}$ factorizes over its latent variables by considering these variables to be independent and each being governed by their own distribution, i.e.,
$q_{\bm{\phi}}(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})$ $=\prod_{d,k=1}^{D,K}q(\theta_{dk})\prod_{k,v=1}^{K,V}q(\beta^{\star}_{kv})\prod_{k,v\in\mathscr{S}}q(\widetilde{\beta}_{kv}).$
Only distributions with support on the positive reals are suitable as variational distributions for document-topic intensities and topic-term intensities. We propose to use Gamma distributions, matching the prior distributions outlined in (2) and (3). We denote shape and rate parameters with the superscript ‘$\mathsf{shp}$’ and ‘$\mathsf{rte}$’, respectively. Moreover, we also employ a scaling by document length $N_{d}=\sum\limits_{v=1}^{V}y_{dv}$ for parameters $\theta_{dk}$. Including the document length $N_{d}$ in this way provided a more stable and quicker model fit and induced a superior classification performance. Altogether, we posit as variational distributions
$q(\theta_{dk})$ $=\Gamma\left(\phi^{\mathsf{shp}}_{\theta_{dk}},N_{d}\cdot\phi^{\mathsf{rte}}_{\theta_{dk}}\right),$ $q(\beta^{\star}_{kv})$ $=\Gamma\left(\phi^{\mathsf{shp}}_{\beta^{\star}_{kv}},\phi^{\mathsf{rte}}_{\beta^{\star}_{kv}}\right),$ $q(\widetilde{\beta}_{kv})$ $=\Gamma\left(\phi^{\mathsf{shp}}_{\widetilde{\beta}_{kv}},\phi^{\mathsf{rte}}_{\widetilde{\beta}_{kv}}\right).$
Hence, we optimize the $\mathsf{ELBO}$ with respect to the set of variational parameters $\bm{\phi}=\{\bm{\phi}^{\mathsf{shp}}_{\theta},\bm{\phi}^{\mathsf{rte}}_{\theta}$, $\bm{\phi}^{\mathsf{shp}}_{\beta^{\star}},\bm{\phi}^{\mathsf{rte}}_{\beta^{\star}},\bm{\phi}^{\mathsf{shp}}_{\widetilde{\beta}},\bm{\phi}^{\mathsf{rte}}_{\widetilde{\beta}}\}\in\bm{\Phi}=\mathbb{R}_{>0}^{2DK}\times\mathbb{R}_{>0}^{2KV}\times\mathbb{R}_{>0}^{2|\mathscr{S}|}$.
We use Black Box Variational Inference (BBVI) with stochastic optimization and follow Ranganath et al., (2014) to form noisy unbiased gradient estimates of the $\mathsf{ELBO}$ with $S$ Monte Carlo samples from the variational distribution,
$\nabla_{\bm{\phi}}\mathsf{ELBO}(\bm{\phi})\approx\frac{1}{S}\sum_{s=1}^{S}\nabla_{\bm{\phi}}\log q_{\bm{\phi}}(\bm{\theta}_{s},\bm{\beta}^{\star}_{s},\widetilde{\bm{\beta}}_{s})\left(\log p(\bm{\theta}_{s},\bm{\beta}^{\star}_{s},\widetilde{\bm{\beta}}_{s},\mathbb{Y})-\log q_{\bm{\phi}}(\bm{\theta}_{s},\bm{\beta^{\star}}_{s},\widetilde{\bm{\beta}}_{s})\right),$ (5)
where $\bm{\theta}_{s},\bm{\beta^{\star}}_{s},\widetilde{\bm{\beta}}_{s}\sim q_{\bm{\phi}}(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})$ are independent samples from the variational distributions. These gradient estimates are used to optimize the $\mathsf{ELBO}$ while the updates $\bm{\phi}$ are determined by the Adam algorithm (Kingma and Ba,, 2015). Reverse-mode automatic differentiation is used to track all sequences of operations and to compute the gradients during the optimization procedure (see Kucukelbir et al.,, 2017). The whole procedure is shown in Algorithm 1 for $S=1$ which is the value for $S$ we use in our implementation. When applying Algorithm 1, we initialize the variational parameter with $\bm{\phi}^{\mathsf{shp}}_{\theta}=1$, $\bm{\phi}^{\mathsf{rte}}_{\theta}=\frac{D}{1000}$, $\bm{\phi}^{\mathsf{shp}}_{\beta^{\star}}=1$, $\bm{\phi}^{\mathsf{rte}}_{\beta^{\star}}=\frac{2D}{1000}$, $\bm{\phi}^{\mathsf{shp}}_{\widetilde{\beta}}=\bm{\phi}^{\mathsf{rte}}_{\widetilde{\beta}}=1$.
Input: DTM $\mathbb{Y}$, number of topics $K$, sets of seed words $\mathscr{V}_{k}$, $k=1,\ldots,K$;
Input: prior parameters $a,b,c,d,e,f$, initial variational parameter $\bm{\phi}$;
Input: number of epochs $E$, batch size $|\mathscr{B}|$, learning rate $\rho$
Output: The last value ^$\bm{\hat{}}{\phi}$ when optimizing $\mathsf{ELBO}(\bm{\phi})$
1
2for epoch $e=1,2,\ldots,E$ do
3       Divide $D$ documents randomly into $B$ batches $\mathscr{B}_{i}$, $i=1,\ldots,B$, $|\mathscr{B}_{i}|\approx|\mathscr{B}|$
4      for i in $1:B$ do
5             for each topic $k\in\{1,\ldots,K\}$ and each word $v\in\{1,\ldots,V\}$ do
6                   Sample $\beta^{\star}_{kv}\sim\Gamma(\phi^{\mathsf{shp}}_{\beta^{\star}_{kv}},\phi^{\mathsf{rte}}_{\beta^{\star}_{kv}})$
7                   if $v\in\mathscr{V}_{k}$ then
8                         Sample $\widetilde{\beta}_{kv}\sim\Gamma(\phi^{\mathsf{shp}}_{\widetilde{\beta}_{kv}},\phi^{\mathsf{rte}}_{\widetilde{\beta}_{kv}})$
9                  else
10                         Set $\widetilde{\beta}_{kv}=0$
11
12            Compute $\bm{\beta}=\bm{\beta^{\star}}+\bm{\widetilde{\beta}}$
13
14            for each document $d$ in batch $\mathscr{B}_{i}$ do
15                   Sample $\theta_{dk}\sim\Gamma(\phi^{\mathsf{shp}}_{\theta_{dk}},N_{d}\cdot\phi^{\mathsf{rte}}_{\theta_{dk}})$
16
17                  for $v\in\{1,\ldots,V\}$ do
18                         Set $\lambda_{dv}=\sum\limits_{k=1}^{K}\theta_{dk}\beta_{kv}$
                         Compute $\log p(y_{dv}\,|\,\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})=\log\mathsf{Pois}(y_{dv}\,|\,\lambda_{dv})$
                          > Log-likelihood
19
20
            Set $\log p(\mathbb{Y}\,|\,\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})=\frac{D}{|\mathscr{B}_{i}|}\sum\limits_{d\in\mathscr{B}_{i}}\sum\limits_{v=1}^{V}\log p(y_{dv}\,|\,\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})$
              > Reconstruction
21
            Compute $\log p(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})$ and $\log q_{\bm{\phi}}(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})$
              > Prior and entropy
22             Compute $\mathsf{ELBO}(\bm{\phi})=\log p(\mathbb{Y}\,|\,\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})+\log p(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})-\log q_{\bm{\phi}}(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})$
23             Compute gradients $\nabla_{\phi}\mathsf{ELBO}(\bm{\phi})$ using automatic differentiation as in (5)
24             Update variational parameter $\bm{\phi}$ with Adam and learning rate $\rho$
25
Algorithm 1 Seeded Poisson factorization algorithm for $S=1$
3.2 Post-processing and final inference
After running the model for a sufficient number of epochs $E$, the final value obtained from the VI optimization $\hat{\bm{\phi}}$ represents the estimate of the variational parameter. We summarize the results by determining point estimates for the parameters of interest based on posterior means derived from the posterior approximations through the variational family. In particular, the posterior mean estimates $\hat{\bm{\theta}},\hat{\bm{\beta}}^{\star},\hat{\widetilde{\bm{\beta}}}$ are obtained by determining the means induced by the variational Gamma distributions. In case of document-topic intensities we obtain posterior means using
$\hat{\theta}_{dk}=\hat{\phi}_{\theta_{dk}}^{\mathsf{shp}}/\hat{\phi}_{\theta_{dk}}^{\mathsf{rte}}.$ (6)
To estimate the topic-term intensities for a topic $k$, we use
$\hat{\beta}_{kv}=\hat{\beta}^{\star}_{kv}+\hat{\widetilde{\beta}}_{kv}=\hat{\phi}_{\beta^{\star}_{kv}}^{\mathsf{shp}}/\hat{\phi}_{\beta^{\star}_{kv}}^{\mathsf{rte}}+\left\{\begin{aligned} &\hat{\phi}_{\widetilde{\beta}_{kv}}^{\mathsf{shp}}/\hat{\phi}_{\widetilde{\beta}_{kv}}^{\mathsf{rte}}&&\text{if}\;(k,v)\in\mathscr{S},\\ &0&&\text{otherwise.}\end{aligned}\right.$ (7)
The final topic assignment is based on a Naive Bayes classifier (see, e.g., Zhang,, 2004), i.e., the document is assigned to the topic where the per-document posterior mean estimate is maximum.
3.3 Computational details
The SPF model is implemented in Python 3.10. It allows Graphics Processing Unit (GPU) support due to its implementation in the TensorFlow environment. The model’s source code is mainly based on TensorFlow 2.15 as well as TensorFlow’s add-on library for probabilistic reasoning, TensorFlow Probability 0.23.0. In its standard implementation, SPF uses a batch size of 1 02410241\,0241 024 documents and trains the model for $150$ epochs. Leveraging TensorFlow’s computational graph and gradient tape functionalities, the implementation enables efficient tracking of operations for automatic differentiation during model training.
The results presented in this paper are compiled locally on a machine with CPU: Intel i5 13600k; GPU: Nvidia RTX 3090; RAM: 32GB DDR5 5600 MHz. Additionally, we employed SPF in an AWS cloud computing environment using an ml.g5dn.xlarge instance [^1] See instance types: https://docs.aws.amazon.com/sagemaker/latest/dg/notebooks-available-instance-types.html. with enhanced GPU support to ensure that the software provided is ready to use in different environments. Our implementation is available as open source software via GitHub. [^2] See https://github.com/BPro2410/Seeded-Poisson-Factorization.
4 Empirical results
4.1 Illustrative application
[…]
When fitting the SPF model, we set the number of topics to correspond to the number of product categories, i.e., $K=6$, and follow the lines of Watanabe and Zhou, (2022) to construct a balanced lexicon of ten frequently occurring seed words for each topic. To generate these seed words, we computed the TF-IDF (Term Frequency-Inverse Document Frequency) values for each word within each category. We select the top 10 words from the TF-IDF matrix as the seed words for each topic. This process is conducted automatically to minimize subjectivity and to ensure that the seed word selection process remains as objective as possible. Table 1 presents a summary of the final sample of documents and the pre-specified seed words per product category. We use the standard specifications for prior and variational parameter values as outlined in Sections 2 and 3 and set the learning rate to $0.1$, the number of epochs to $150$ and the batch size to 1 02410241\,0241 024.
Product category Topic Count Seed words Toys games Toys 8 092 toy, game, play, fun, old, son, year, loves, kids, daughter Health personal care Health 6 938 product, like, razor, shave, time, day, shaver, better, work, years Beauty products Beauty 4 072 hair, skin, product, color, scent, smell, used, dry, using, products Baby products Baby 4 635 baby, seat, diaper, diapers, stroller, bottles, son, pump, gate, months Pet supplies Pets 3 792 dog, cat, litter, cats, dogs, food, box, collar, water, pet Grocery gourmet food Grocery 2 471 tea, taste, flavor, coffee, sauce, chocolate, sugar, eat, sweet, delicious
Table 1: Overview of the final sample: document counts and seed words by product category.
4.1.1 Topic interpretability
We examine the interpretability of the topics inferred by SPF based on the approximate posterior mean topic-term intensities $\hat{\bm{\beta}}_{k}$ for each topic $k$. Table 2 presents the top 14 terms with the highest approximate mean intensities per topic, after removing stopword-like terms that provide no contextual information. Bold terms represent seed words. The mean term intensity (provided in parentheses) is calculated as the sum of the seeded term intensity and the unseeded term intensity, as defined in Equation 7. These high-intensity words per topic enable the characterization of the topic as well as assess how influential the seed words were.
Clearly the pre-defined seed words exhibit a strong presence among the most pertinent terms for all topics. However, the number of seed words included in the top 14 words with highest intensity varies across topics. For topic ‘Toys‘, all 10 seed words are also included in the list of 14 most pertinent terms for this topic. This number decreases to six for ‘Health’, eight for ‘Beauty’, seven for ‘Baby’ and eight for ‘Pets’. The lowest number of seed words are included in the list of 14 most pertinent terms for the topic ‘Grocery’ where only four out of the ten seed words are listed.
Table 2 reveals that the model effectively prioritizes not only the explicitly defined seed words but also identifies and assigns significant weight to relevant additional terms that are not pre-specified as seed words. E.g., for topic ‘Health’ the seed word ‘day’ was specified and also ‘days’ is included among the 14 most pertinent terms. For topic ‘Baby’ not only the seed word ‘son’ is included but also ‘daughter’. Inspecting the ‘Baby’ topic further by also assessing additional terms with high intensity indicates that SPF did not only assign high relevance to expected seed words such as ‘seat’ (14.88) and ‘son’ (12.07), but also recognized terms like ‘bed’ (6.71) and ‘sleep’ (6.31) as highly pertinent to the topic. These terms align with the product subcategory ‘sleep positioners’, which falls under the broader ‘Baby’ category, demonstrating the model’s nuanced understanding of topic content and its ability to discern contextually important terms.
The topic ‘Grocery’ fails to capture most of the seed words among the 14 most pertinent terms. However, the terms with high intensity suggest that this topic captured an additional aspect in customer reviews which relates not to the product but the purchase process. For example, terms like ‘store’ ($7.34$) and ‘shipping’ ($7.20$) have a high prevalence in the ‘Grocery’ topic.
Toys Health Beauty Baby Pets Grocery toy (39.67) product (25.40) hair (48.66) baby (32.00) dog (24.57) amazon (19.30) old (30.27) time (17.00) product (30.67) use (20.89) water (16.68) like (18.62) game (22.08) work (14.58) like (24.75) seat (14.88) cat (15.62) product (16.65) play (21.54) years (13.01) use (22.91) easy (14.60) box (14.19) taste (13.70) year (20.48) day (12.70) skin (22.41) little (12.92) product (12.05) tea (13.16) fun (19.22) used (11.64) really (13.63) son (12.07) dogs (10.55) price (10.24) loves (18.70) good (9.92) color (12.06) old (11.57) cats (10.53) flavor (9.80) great (18.23) works (7.48) smell (11.93) months (11.25) litter (9.94) buy (7.66) like (17.62) days (7.24) dry (10.64) fit (10.13) small (9.31) store (7.34) little (17.09) batteries (7.19) time (10.22) car (9.92) time (8.78) shipping (7.20) son (16.68) battery (6.55) good (9.90) daughter (9.08) little (8.29) order (7.10) daughter (14.85) pain (6.43) products (8.83) diaper (8.01) plastic (8.05) eat (6.02) bought (12.56) razor (6.29) face (8.79) bottles (7.24) clean (7.46) food (6.00) kids (12.52) reviews (6.29) scent (8.44) pump (6.97) food (7.41) protein (5.44)
Table 2: High-intensity words per topic. Mean intensities are shown in brackets. Bold words are seed words.
Table 2 also illustrates that SPF is able to assign distinctly different intensities to the seed words as well as other terms with high intensity within their respective topics. E.g., within the ‘Pets’ category, ‘dog’ ($24.57$) plays a dominant role, whereas other seed words, like ‘food’ ($7.41$), display a markedly lower mean intensity. This contrast highlights the ability of SPF to estimate the uneven influence of seed words in defining a topic. This property of the SPF topic model is important to also mitigate the risk of potential mis-specifications that may arise due to incomplete domain knowledge during the initial selection of seed words. To empirically assess the influence of mis-specified seed words, we conducted an additional analysis evaluating the performance of SPF when an inappropriate seed word is assigned to a topic. Specifically, we fitted the SPF model with ‘dog’ as a seed word for the ‘Beauty’ topic. The results indicate that SPF effectively recognizes that the term contributes minimal to no informational value in this context. In particular, the inferred variational distribution was $\widetilde{\beta}_{beauty,dog}\sim\Gamma(0.25,5.66)$. These findings underscore SPF’s ability to adaptively assign importance to seed words, thereby reducing the impact of initial specification errors.
4.1.2 Predictive performance
Next, we measure the predictive performance of SPF based on approximate posterior means of the document-topic intensities, where each document vector is a $K$-dimensional vector of approximate mean intensities $\hat{\bm{\theta}}_{d}\in\mathbb{R}_{>0}^{K}$, see Equation 6. According to the Naive Bayes classifier, the topic with the highest approximate mean intensity in $\hat{\bm{\theta}}_{d}$ is assigned as the predicted topic for document $d$. We assess the predictive performance separately for each topic based on precision (correctly assigned to the topic among all assigned to the topic), recall (correctly assigned to the topic among all from the topic) and F1-score (harmonic mean of precision and recall) as well as on aggregate with equal category weights (macro avg) and taking the empirical category imbalance into account (weighted avg). We also provide information on assignment certainty, presenting insights into the model’s confidence in its predictions per category, by determining the fraction of document-topic intensity the topic has the document is assigned to. Specifically, we show the average assignment certainty for true positive (TP) and false positive (FP) predictions.
True label Precision Recall F1-score Assignment certainty Count TP FP Toys 0.92 0.82 0.87 0.68 0.51 8 092 Health 0.75 0.46 0.57 0.64 0.60 6 938 Beauty 0.68 0.79 0.73 0.71 0.54 4 072 Baby 0.71 0.78 0.74 0.70 0.54 4 635 Pets 0.61 0.76 0.74 0.65 0.52 3 792 Grocery 0.51 0.94 0.66 0.78 0.54 2 471 Macro avg 0.71 0.76 0.72 30 000 Weighted avg 0.75 0.73 0.73 30 000
Table 3: Predictive performance of the SPF topic model on Amazon customer feedback, including assignment certainties of TP (true positive) and FP (false positive) predictions. The overall accuracy is 0.73.
The classification performance of the SPF topic model is summarized in Table 3. Clearly, SPF provides excellent classification performance in categorizing Amazon customer feedback across all six product categories, despite slight differences across the categories. The overall accuracy of the model is 0.73, which is consistent with the weighted average F1-score (0.73), accounting for the varying sample sizes across categories. The macro average F1-score (0.72) is slightly lower, reflecting the imbalanced performance among categories.
The model exhibits varying performance across categories. For instance, the highest F1-score is observed for the ‘Toys’ category ($0.87$), reflecting both high precision ($0.92$) and recall ($0.82$). This suggests strong alignment between predicted and true labels. Conversely, the ‘Grocery’ category achieves the lowest F1-score ($0.66$), driven by a significant imbalance between precision ($0.51$) and recall ($0.94$). This discrepancy indicates a tendency to over-predict the ‘Grocery’ category, resulting in higher recall at the cost of precision. The over-prediction of the ‘Grocery’ topic is also reflected in the highest TP assignment certainty score of $0.78$. By contrast, the ‘Health’ category shows an inverted pattern, with a high precision ($0.75$) but a much lower recall ($0.46$), indicating under-representation in predictions. In the ‘Health’ category, the model exhibits the lowest TP assignment certainty at $0.64$, indicating that it is on average less confident in the assignment compared to all other categories when correctly assigning a document. At the same time, the FP assignment certainty is the highest among all categories at $0.60$. This combination of low confidence in true positives and high confidence in false positives highlights the model’s particular struggle with distinguishing health-related feedback, emphasizing the need for further refinement in this category. Table 2 shows that the topic-term intensities of seed words for the ‘Grocery’ and ‘Health’ category are in general not as strong as the ones for the other categories. This likely contributes to the lower classification performance observed for these categories. The correct specification of seed words appears to be a crucial factor in achieving high classification performance, as demonstrated by the results in the ‘Toys’ category. This highlights the model’s particularly strong performance in categories with distinct linguistic characteristics. However, categories like ‘Grocery’ and ‘Health’ reveal areas where the model might benefit from further refinement, such as enhanced domain-specific seed words or adjustments to address label imbalance. Also including an additional unseeded topic which captures feedbacks discussing the purchasing process rather than the product could improve the categorization of the feedback items. The overall accuracy and balanced macro and weighted averages suggest a generally robust model. Further refinement could enhance performance in underperforming categories.
4.2 Comparison to existing methods and scalability
To evaluate the predictive performance and the computational efficiency of the SPF topic model in comparison with other guided topic models, we also fit KeyATM (Eshima et al.,, 2024) and SeededLDA (Watanabe and Zhou,, 2022) to the Amazon corpus. We investigate in particular how the predictive performance and run-times change with the number of documents $D$ in the corpus, varying $D$ from $1k$, over $5k$ and $10k$ to $30k$. The evaluation criteria include the run-time (in minutes) as well as the predictive performance metrics accuracy, precision, recall, and F1-score. These computational experiments were conducted using the hardware setup described in Section 3.3.
For all models, we set the number of topics to $K=6$ and limited the input data to customer feedback only, excluding any additional metadata. Each model was trained using the same set of seed words and the default values for model and prior specifications suggested / implemented in the software packages. Using the same number of MCMC iterations for KeyATM and SeededLDA as well as number of epochs for model fitting regardless of the number of documents $D$ did lead to poor predictive performance results in case only very few documents were included in the corpus. We thus increased the number of MCMC iterations / number of epochs for $D$ equal to $5k$ and $1k$. In particular we used 1 500 MCMC iterations and 150 epochs for $D\in\{10k,30k\}$ and doubled this number to 3 000 MCMC iterations and 300 epochs for $D=5k$ and tripled the number to 4 500 MCMC iterations and 450 epochs for $D=1k$.
Table 4 provides the results for this comparison. The run-time comparison clearly shows that regardless of corpus size SPF always has the shortest run-times compared to SeededLDA and finally KeyATM. The difference in run-time increases with the corpus size. While for a corpus of size $1k$ documents, the run-times of KeyATM and SeededLDA are only approximately 1.5 times the run-time of SPF, the run-times increase by a factor of 3 to 5 times for a corpus of size $30k$ documents.
In terms of accuracy, SPF demonstrates superior or on par performance compared to the other methods across all corpora sizes. For the largest corpus ($D=30k$), SPF achieves an accuracy of $0.73$, equivalent to that of KeyATM and higher than SeededLDA’s $0.65$. As the corpus size decreases, SPF maintains a higher accuracy, notably outperforming the competing methods for smaller corpora. In particular, SPF achieves high accuracy up to a corpus size of $5k$ with a considerable drop in performance only observed in case $1k$ documents are included in the corpus. Also SeededLDA maintains a similar – although at a lower level – accuracy regardless of corpus size with only a slight reduction in case of a small corpus size. KeyATM is most severely effected by a decrease in corpus size with the accuracy values dropping from 0.73 to 0.29. SPF also outperforms the other methods with respect to the other predictive performance criteria such as precision, recall and F1-score.
$30k$ documents $10k$ documents SPF KeyATM SeededLDA SPF KeyATM SeededLDA Time (min:sec) 1:07 5:27 3:35 0:19 1:41 1:04 Accuracy 0.73 0.73 0.65 0.72 0.57 0.63 Precision 0.71 0.73 0.63 0.71 0.72 0.62 Recall 0.76 0.72 0.68 0.75 0.52 0.66 F1-score 0.72 0.72 0.65 0.71 0.55 0.63 $5k$ documents $1k$ documents SPF KeyATM SeededLDA SPF KeyATM SeededLDA Time (min:sec) 0:09 0:44 0:34 0:06 0:09 0:09 Accuracy 0.70 0.44 0.62 0.63 0.29 0.58 Precision 0.68 0.70 0.61 0.61 0.50 0.57 Recall 0.73 0.38 0.64 0.67 0.21 0.59 F1-score 0.69 0.39 0.62 0.62 0.15 0.57
Table 4: SPF vs. existing methods with different corpora sizes.
[…]
Figure 3: Processing time for the bootstrap experiment.
4.3 Robustness checks
We systematically vary key parameters of the specification and estimation of the SPF topic model to evaluate how the performance changes under different conditions. Table 5 presents the resulting 13 scenarios and the corresponding predictive performance results obtained when fitting SPF under these scenarios.
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11) (12) (13) Setting Number of documents 30k 30k 30k 30k 30k 30k 30k 30k 30k 30k 30k 1k 1k Seeded topics 6 6 6 6 6 6 6 5 6 6 6 6 6 Unseeded topics 0 0 0 0 0 0 0 1 1 0 0 0 0 $\widetilde{\beta}$ shape prior parameter 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 0.3 2.0 0.3 2.0 Seed words per topic 10 5 10 10 10 10 10 10 10 10 10 10 10 Batch size 1 024 1 024 1 024 1 024 512 512 512 1 024 1 024 1 024 1 024 1 024 1 024 Learning rate 0.1 0.1 0.01 0.01 0.1 0.01 0.01 0.1 0.1 0.1 0.1 0.1 0.1 Epochs 150 150 150 300 150 150 300 150 150 150 150 300 300 Performance metrics Accuracy 0.73 0.70 0.50 0.66 0.73 0.64 0.71 0.73 0.67 0.73 0.73 0.62 0.64 Precision 0.71 0.68 0.58 0.65 0.71 0.64 0.69 0.71 0.65 0.71 0.72 0.60 0.62 Recall 0.76 0.74 0.55 0.70 0.76 0.68 0.74 0.76 0.60 0.76 0.76 0.65 0.68 F1-score 0.72 0.69 0.51 0.65 0.72 0.63 0.69 0.72 0.61 0.72 0.72 0.61 0.63
Table 5: Robustness checks. Analysis of 13 different scenarios with varying settings regarding the data, model specification and estimation. The setting for the benchmark model is shown in Scenario (1). Changes in the settings, compared to the base scenario, are indicated by a gray background.
Effect of the number of seed words.
We examine the impact of reducing domain knowledge on classification performance by reducing the number of seed words per topic from 10 to 5 (Scenario 2). As expected, decreasing the amount of seed words results in lower predictive performance. However, halving the number of seed words did result only in a slight decrease in predictive performance, underscoring the importance of being able to select at least a few meaningful seed words to characterize topics in order to achieve superior results using the seeded approach when fitting topic models.
Effect of learning rate, epochs and batch size.
In Scenarios 3, 4, 6 and 7 the learning rate is reduced to 0.01 from 0.1 in the base scenario. Results indicate that in particular lowering the learning rate increases the number of epochs required for the negative $\mathsf{ELBO}$ to converge, suggesting slower optimization (see Scenario 4).
For the scenarios considered, increasing the number of epochs (Scenarios 4, 7, 12, 13) does not improve the performance. However, a higher number of epochs was in particular used when the simultaneous change of another setting induced the need for more epochs, such as the reduction of the learning rate (Scenarios 4 and 7) or a lower number of documents (Scenarios 12 and 13).
Reducing the batch size (Scenarios 5, 6, 7) seems to have hardly any impact. Similar results are obtained in particular for Scenario 5 where this is the only setting change. Scenario 7 suggests that lowering the learning rate, reducing the batch size and increasing the number of epochs yields good results, highlighting the interplay between these hyperparameters.
Effect of varying the number of topics $K$.
We also examine the effect of altering the number of topics $K$ in two different ways. First, we remove the a-priori information for the ‘Grocery’ topic, estimating the model with five seeded topics and one unseeded topic (Scenario 8). Second, we add an additional unseeded topic to the six seeded product categories (Scenario 9).
Dropping the seed words for one topic but otherwise keeping the number of topics (Scenario 8) results in them same good performance than in the base scenario. In the case where ‘Grocery’ topic is excluded and an unseeded topic is added, SPF allocates 4 82648264\,8264 826 customer reviews to the unseeded category, accurately identifying 2 33123312\,3312 331 out of 2 47124712\,4712 471 instances as belonging to the ‘Grocery’ category. Inspecting the words with the highest topic-term intensities for the unseeded topic indicates that this topic effectively captures the ‘Grocery’ topic. High-intensity words emerging in this setting are ‘taste’ (12.84), ‘tea’ (12.13), and ‘flavor’ (8.78).
In Scenario 9, adding an additional unseeded topic leads to reduced model performance ($0.67$ compared to $0.73$ in the baseline scenario), as the fixed number of six product categories means that assigning a customer review to the unseeded topic constitutes a mis-classification in this context. However, an analysis of the topic-term distributions for the unseeded topic reveals that SPF assigns reviews to the unseeded topic when customers primarily discuss the purchasing process rather than specific product characteristics. To give an example, terms such as ‘time’, ‘shipping’, ‘store’, ‘order’ and ‘online’ exhibit high intensities within the unseeded topic. SPF was hence capable to identify the additional latent topic present in customer reviews which relates to the purchasing and delivery experiences, which, while not tied to specific product categories, is also of significant business relevance. Overall these findings illustrate SPF’s robust capability to discern meaningful latent topics even in the absence of comprehensive domain-specific seeding.
Effect of the shape parameter $c$ on $\widetilde{\beta}$.
[…]
5 Discussion
Traditional topic models often struggle to align the latent topics they derive with pre-specified concepts of interest (see, e.g., Eshima et al.,, 2024). To address these limitations, we extend the Poisson factorization topic model with a seeded approach. The seeded approach guides the inference of topics, avoiding the need for manual labeling, but also enables the use of topic models for text classification where labeled text data are not available but the classes for categorization are readily characterized by a set of relevant words. Seeding modifies the prior distribution of the topic-term distributions by assigning higher rates a-priori to the relevant words associated with their respective topics.
Our empirical findings demonstrate that integrating domain knowledge into the model specification significantly enhances the capability of topic models to extract meaningful topic-term intensities, thereby improving the interpretability and understanding of topics. Additionally, by applying a Naive Bayes classifier based on the fitted document-topic distributions, we are able to classify documents in an automatic way. Experiments on a pre-labeled dataset reveal that the SPF topic model not only improves topic interpretability but also achieves superior predictive performance compared to alternative seeded topic modeling approaches. By combining the computational efficiency of VI techniques with the prior knowledge of domain experts in a PF framework, SPF enables a robust and effective system for document classification. This synergy improves the overall quality and utility of the classification process, making it more reliable and actionable for a wide range of applications.
Acknowledgments
The authors gratefully acknowledge support from the Jubiläumsfonds of the Oesterreichische Nationalbank (OeNB, grant no. 18718).